# PAPI-Reg: Patch-to-Pixel Solution for Efficient Cross-Modal Registration between LiDAR Point Cloud and Camera Image

### v1

@@ Abstract
The primary requirement for cross-modal data fusion is the precise alignment of data from different sensors. However, the calibration between LiDAR point clouds and camera images is typically time-consuming and needs external calibration board or specific environmental features. Cross-modal registration effectively solves this problem by aligning the data directly without requiring external calibration. However, due to the domain gap between the point cloud and the image, existing methods rarely achieve satisfactory registration accuracy while maintaining real-time performance. To address this issue, we propose a framework that projects point clouds into several 2D representations for matching with camera images, which not only leverages the geometric characteristic of LiDAR point clouds more effectively but also bridge the domain gap between the point cloud and image. Moreover, to tackle the challenges of cross modal differences and the limited overlap between LiDAR point clouds and images in the image matching task, we introduce a multi-scale feature extraction network to effectively extract features from both camera images and the projection maps of LiDAR point cloud. Additionally, we propose a patch-to-pixel matching network to provide more effective supervision and achieve higher accuracy. We validate the performance of our model through experiments on the KITTI and nuScenes datasets. Our network achieves real-time performance and extremely high registration accuracy. On the KITTI dataset, our model achieves a registration accuracy rate of over 99%.
1 Introduction
@@
In autonomous driving and robot systems, the fusion between data collected from different sensors now plays an important role in many downstream tasks, such as object detection [1, 6, 15] and segmentation [5, 33, 31]. We can achieve better accuracy and efficiency by using these fused data, thereby ensuring safety and robustness in complex environments. Among these methods, the most common process is to perform a fusion of the LiDAR (solid-state LiDAR or rotating LiDAR, etc.) point cloud and the camera (high-definition cameras or infrared cameras, etc.) image.
@@
Figure 1: The problem about the domain gap in the cross-modal registration task. Due to the dimensional differences between point clouds and images, the matching failure rate between them is high, which in turn reduces the registration accuracy.
@@
However, the primary condition of a successful fusion is an accurate alignment between these data. A commonly used method to deal with the alignment task between sensors is to calculate the external parameters between sensors using a checkerboard [14], which we call off-line calibration. We can make a high-accuracy alignment between sensors by using these types of methods and unavoidable time costs. In addition, some online calibration methods [27, 13, 36] are proposed to correct external parameter errors caused by bumps or vibrations. These online calibration methods can modify the deviations within a limited error range. However, they will lose effectiveness when the errors are too large. Although these methods can achieve high-accuracy alignment, these complicated calibration processes make data fusion preprocessing time-consuming and inefficient. Whenever the deployment of sensors has changed, these processes have to be repeated.
@@
Why not propose a method to bypass these calibration processes and save manpower and resources by directly doing a registration between the collected data? Cross-modal registration methods have emerged to address this challenge. Recent cross-modal registration methods [9, 21] can solve this cross-modal data alignment task by using specific deep learning networks. However, the cross-modal registration process typically involves comparing the similarity of the high-dimensional features between point clouds and images to establish correspondences, and the domain gap between 3D and 2D data always influence this process, as we showed in Fig. 1. Specifically, there is a dimensional and informational difference between the 3D spatial data of the point cloud and the 2D texture information of the image. Although some existing methods attempt to decline these influence by exchanging feature information [24] or using feature extraction networks with more similar structures [39], this domain gap still affect the accuracy of the registration, especially leading to a high mismatching rate.
@@
Figure 2: The challenge we faced in the matching task between projection map and camera image. The cross-modal difference between the camera image and the projection map will make it difficult to the feature matching. Additionally, the scarcity of positive samples caused by the limited overlapping region will lead to the network’s supervision problem.
@@
Besides, the commonly used point cloud feature extraction network in the cross-modal registration task (e.g., PointNet++ [23], sparse convolution network[7]) always treats the LiDAR point cloud simply as a series of unordered points or voxels, which fails to effectively use the spatial and geometric characteristics of the point cloud. Therefore, to make better use of the scanning characteristics of LiDAR to bridge the domain gap, we introduce a method to project point clouds into 2D representations: a range map and a reflectance map. Although using these projection methods is not new in other downstream tasks involving LiDAR point clouds, they have not been fully leveraged in cross-modal registration methods. Moreover, mapping point clouds to pixels in 2D representations offers a data format that implicitly contains more geometric information, making it more effective and accurate than using unordered sparse point clouds. This projection approach will enable more effective feature extraction and bridge the domain gap difference between point clouds and images. Furthermore, we combined the use of the two projection maps containing different types of information to enhance the system’s robustness for registration across various scenarios.
@@
After the projection, we transform the cross-modal registration task to an image registration task between the camera image and the projection maps. Common image registration methods are generally accomplished through key point matching. Due to the cross-modal differences between the projected 2D map and the camera image, traditional key point registration methods cannot be directly applied to this task, as we showed in Fig. 2. Furthermore, the overlapping region between the 2D map and the camera image is often limited in practical scenarios, therefore the proportion of truly matched point pairs becomes extremely small, leading to positive pair scarcity. For existing deep learning-based image matching methods, this scarcity can cause supervision difficulties, making the network hard to converge and increasing the number of outliers, which in turn raises the likelihood of registration failures.
@@
To address these challenges, we designed a multi-scale feature extraction network to mitigate the cross-modal differences between images and projection maps, effectively extracting features. Additionally, we introduced a patch-to-pixel feature matching framework. Specifically, patch-level searching addresses the supervision and success rate issues caused by positive sample scarcity, while pixel-level matching enhances registration accuracy. Our experiment part has proved that our network design achieves high accuracy while maintaining low computational complexity. In conclusion, our contributions are as follows.
@@
    We propose a robust registration framework that leverages range map and reflectance map features for cross-modal registration between LiDAR point clouds and camera images, which can achieve a registration accuracy rate of over 99% in the KITTI dataset.
@@
    We introduce a patch-to-pixel feature matching framework, which not only resolves the supervision problems that arise from the limited overlap between the projection map and the camera image but also enhances registration accuracy.
2 Related Work
2.1 Off-line Calibration Methods
Traditional offline extrinsic calibration methods [14, 40] are typically conducted before system operation, using manually placed prominent markers to identify 2D-3D correspondences and then estimate the external parameters between sensors using the least-square method [38]. These markers, such as checkerboard, generally contain distinct features on points [12], lines [2], or planes [37]. After establishing the correspondence between 2D and 3D feature points [29]. Such methods are performed before system deployment and can achieve high calibration accuracy. However, the marker placement and calibration process often require significant manpower and resources. If the sensor system’s relative poses are actively changed or vibrations or shocks during operation cause sensor position shifts, the calibration needs to be repeated, which is time-consuming and labor-intensive.
2.2 Online Calibration Methods
Subsequently, many online calibration methods were proposed to optimize minor errors that occur during system operation, such as those caused by vibrations, after initial offline calibration. Existing approaches include using MLPs to estimate deviations [36, 13, 19], keypoint-based methods[35], and techniques that leverage objects [28] or segmentation [42, 32]. These real-time optimization techniques improve data alignment and enhance data fusion, leading to better performance in downstream tasks. However, these methods rely on the initial extrinsic parameters obtained from prior calibration and lose effectiveness when there are significant deviations between the actual and initial parameters. Moreover, deploying online calibration on top of complex offline calibration processes further increases costs.
2.3 Cross-modal registration Methods
Cross-modal registration methods are often used in visual localization tasks [3, 25, 41], where the goal is to estimate the camera extrinsic parameters in the world coordinate system defined by large-scale point clouds. Traditionally, these methods are based on 2D visual descriptors [8, 26] to achieve alignment. Later advancements have introduced cross-dimensional descriptors [9, 21, 30] designed for direct cross-modal matching. These methods generally involve extracting features from different modalities to create high-dimensional representations, using feature similarity in this space to establish correspondences. Subsequently, transformation matrices are estimated using techniques like EPnP [16].
@@ 2.3 Cross-modal registration Methods
Figure 3: Visualizations of the projection map, (a) Projected map by spherical projection. The red area represents the empty holes. (b) Projected map from modified projection method using LaserID. The red area represents the empty holes. (c) The detailed process of modified projection method using LaserID.
3 Proposed Method
3.1 Problem Statement
@@
3.2 Data Pre-processing
@@
To get a better use of the geometric characteristics of LiDAR point cloud, and mitigate the impact of domain gap between the point cloud and the image, we project the original 3D point cloud into 2D representations by a specific projection method. Currently, there are various methods for such projection, such as spherical projection [6], cylindrical projection [20], and bird’s-eye view (BEV) projection [34]. Among them, the spherical projection method can make good use of the characteristics of lateral scanning of LiDAR. However, according to the research by Wu et al. [34], direct spherical projection may result in numerous empty holes in the projected map, due to slight misalignments between the coordinate systems of individual lasers and the overall LiDAR system, as shown in the red area of Figure 3(a). These empty areas could negatively impact subsequent registration performance.
To address this issue, we use a modified the spherical projection by associating the vertical axis of the projection map with the LaserID of the LiDAR point cloud instead of the elevation angle, which can reduce the empty holes effectively, as illustrated in Figure 3(b). By using the radius in the spherical coordinate system as the information projected onto the map, we obtain the range map $\mathbf{I}_{Ra}\in\mathbb{N}^{H_{R}\times W_{R}}$. Similarly, projecting the reflectance intensity yields the reflectance map $\mathbf{I}_{Rf}\in\mathbb{N}^{H_{R}\times W_{R}}$. These maps, along with the camera image $\mathbf{I}_{C}\in\mathbb{N}^{H\times W\times 3}$, are fed into the subsequent network. Some newer and more advanced multi-line LiDAR sensors may not experience these misalignment issues. In such cases, spherical or cylindrical projection methods can be used directly.
To explain this projection process vividly, we take the reflectance map $\mathbf{I}_{Rf}$ as an example, as shown in Figure 3(c). For an $N$-line LiDAR, because the horizontal axis of the projection map corresponds to the azimuth angle $\phi$ in the spherical coordinate system and the vertical axis corresponds to the LaserID, we can assume that we have two data points at the same azimuth angle $\phi$: point A in the first line (LaserID = 0) with spherical coordinates $(r_{A},\theta_{A},\phi)$ and point B in the last line (LaserID = $N$) with coordinates $(r_{B},\theta_{B},\phi)$, where, $r_{A}$ and $r_{B}$ are the spherical radii of points A and B, and $\theta_{A}$ and $\theta_{B}$ are their pitch angles. After this LaserID-based projection, they are mapped to the same column $u_{\phi}$, with Pixel A at $(u_{\phi},0)$ in the first row and Pixel B at $(u_{\phi},N)$ in the last row of the reflectance map.
3.3 Feature Extraction Network
Figure 5: The detailed architecture of the patch-to-pixel matching network.
@@ 3.3 Feature Extraction Network
We employ convolutional neural networks (CNNs) to process the camera image, range map, and reflectance map. The camera encoder extracts features from $\mathbf{I}_{C}$, producing a high-dimensional feature at 1/32 of the original image scale. This small-scale, high-dimensional feature is then passed to the camera decoder. In the camera decoder network, the feature scale is gradually upsampled. First, we obtain the patch-level feature at 1/4 of the original image scale, $\mathbf{f}_{C}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times D_{\text{patch}}}$, where $D_{patch}$ is the number of patch-level feature channels. Then after two layers of transposed convolution to achieve pixel-level features $\mathbf{F}_{C}\in\mathbb{R}^{H\times W\times D_{\text{pixel}}}$ at the original image resolution, where $D_{pixel}$ is the number of pixel-level feature channels.
@@
3.4 Patch-to-Pixel Matching Network
@@
The structure of our matching network is shown in Figure 5. Initially, we input the patch-level features $\mathbf{f}_{C}$ and $\mathbf{f}_{L}$, extracted by the feature extraction network, into a matching module to perform patch-level matching. The matching module is similar to the optimization network structure in Lightglue [18]. First, we compute the score matrix $\mathbf{S}$ between $\mathbf{f}_{C}$ and $\mathbf{f}_{L}$. Specifically, these features are flattened and then passed through a linear layer before performing matrix multiplication.

### v2

# EEPNet-V2: Patch-to-Pixel Solution for Efficient Cross-Modal Registration between LiDAR Point Cloud and Camera Image

@@ Abstract
The primary requirement for cross-modal data fusion is the precise alignment of data from different sensors. However, the calibration between LiDAR point clouds and camera images is typically time-consuming and needs external calibration board or specific environmental features. Cross-modal registration effectively solves this problem by aligning the data directly without requiring external calibration. However, due to the domain gap between the point cloud and the image, existing methods rarely achieve satisfactory registration accuracy while maintaining real-time performance. To address this issue, we propose a framework that projects point clouds into several 2D representations for matching with camera images, which not only leverages the geometric characteristic of LiDAR point clouds effectively but also bridge the domain gap between the point cloud and image. Moreover, to tackle the challenges of cross modal differences and the limited overlap between LiDAR point clouds and images in the image matching task, we introduce a multi-scale feature extraction network to effectively extract features from both camera images and the projection maps of LiDAR point cloud. Additionally, we propose a patch-to-pixel matching network to provide more effective supervision and achieve high accuracy. We validate the performance of our model through experiments on the KITTI and nuScenes datasets. Experimental results demonstrate the the proposed method achieves real-time performance and extremely high registration accuracy. Specifically, on the KITTI dataset, our model achieves a registration accuracy rate of over 99%. Our code is released at: https://github.com/ESRSchao/EEPNet-V2.
Index Terms:
LiDAR-camera registration, cross-modal data fusion, deep learning, point cloud projection.
I Introduction
@@
In autonomous driving and robot systems, the fusion between data collected from different sensors now plays an important role in many downstream tasks, such as object detection [1; 2; 3] and segmentation [4; 5; 6]. Better accuracy and efficiency can be achieved by using these fused data, thereby ensuring safety and robustness in complex environments. Among the data collected from different sensors, the most common process is to perform a fusion of the LiDAR (solid-state LiDAR or rotating LiDAR, etc.) point cloud and the camera (high-definition cameras or infrared cameras, etc.) image.
@@
Figure 1: The domain gap in the cross-modal registration task. Due to the dimensional differences between point clouds and images, the matching failure rate is high, which in turn reduces the registration accuracy. The final graph is the projection result using a wrongly estimated extrinsic matrix
@@
However, the primary condition of a successful fusion is an accurate alignment between these data. A commonly used method to deal with the alignment task between sensors is to calculate the external parameters between sensors using a checkerboard [7], which we call off-line calibration. High-accuracy alignment between sensors by using these types of methods and unavoidable time costs. In addition, some online calibration methods [8; 9; 10] are proposed to correct external parameter errors caused by bumps or vibrations. These online calibration methods can modify the deviations within a limited error range. However, they will lose effectiveness when the errors are too large. Although these methods can achieve high-accuracy alignment, these complicated calibration processes make data fusion preprocessing time-consuming and inefficient. Whenever the deployment of sensors has changed, these calibration processes have to be repeated.
@@
To save manpower and resources, a straightforward way is to bypass these calibration processes by directly registering the data collected data from different sensors. Therefore cross-modal registration methods have emerged to address this challenge. Recent cross-modal registration methods [11; 12] can solve this alignment task by using specific deep learning networks. However, the cross-modal registration process typically involves comparing the similarity of the high-dimensional features between point clouds and images to establish correspondences, and the domain gap between 3D and 2D data always influence this process, as shown in Fig. 1. Specifically, there is a dimensional and informational difference between the 3D spatial data of the point cloud and the 2D texture information of the image. Although some existing methods attempt to decline these influence by exchanging feature information [13] or using feature extraction networks with more similar structures [14], this domain gap still affect the accuracy of the registration, especially leading to a high mismatching rate.
@@
Figure 2: The challenge we faced in the matching task between projection map and camera image. The cross-modal difference between the camera image and the projection map will make it difficult for the feature matching. Additionally, the scarcity of positive samples caused by the limited overlapping region will lead to the network’s supervision problem.
@@
Besides, the commonly used point cloud feature extraction network in the cross-modal registration task (e.g., PointNet++ [15], sparse convolution network[16]) always treats the LiDAR point cloud simply as a series of unordered points or voxels, which fails to effectively use the spatial and geometric characteristics of the point cloud. Therefore, to make better use of the scanning characteristics of LiDAR to bridge the domain gap, we introduce a method to project point clouds onto 2D representations: a range map and a reflectance map. Although using these projection methods is not new in other downstream tasks involving LiDAR point clouds, they have not been fully leveraged in cross-modal registration methods. Moreover, mapping point clouds to pixels in 2D representations offers a data format that implicitly contains more geometric information, making it more effective and accurate than using unordered sparse point clouds. This projection approach will enable more effective feature extraction and bridge the domain gap difference between point clouds and images. Furthermore, we combined the use of the two projection maps containing different types of information to enhance the system’s robustness for registration across various scenarios.
@@
After the projection, we transform the cross-modal registration task to an image registration task between the camera image and the projection maps. Common image registration methods are generally accomplished through key point matching. Due to the cross-modal differences between the projected 2D map and the camera image, traditional key point registration methods cannot be directly applied to this task, as the callenge 1 we showed in Fig. 2.
Furthermore, we previously proposed EEPNet [17], which leverages point cloud projection to mitigate domain discrepancies between data of different dimensions. By matching edge pixels between the projected maps and camera images, it effectively reduces computational complexity. This method achieves highly accurate registration results on the KITTI dataset. Notably, due to the projection strategy and edge information utilization, the inference time is remarkably short, reaching as low as 15ms per frame. However, a limitation of this approach is that its training strategy appears effective only for dense point clouds, such as those in the KITTI dataset. When applied to sparse point clouds (e.g., nuScenes), the training loss fails to converge. Our analysis reveals that this issue stems from an insufficient proportion of matched edge pixels relative to the total pixels, which restricts EEPNet’s robustness and limits its applicability to broader scenarios, as shown in the callenge 2 of Fig. 2.
@@
To address these challenges, we design a multi-scale feature extraction network to mitigate the cross-modal differences between images and projection maps, effectively extracting features. Additionally, we introduce a patch-to-pixel feature matching framework. Specifically, patch-level searching addresses the supervision and success rate issues caused by positive sample scarcity, while pixel-level matching enhances registration accuracy. Experimental results show that the proposed method achieves high accuracy while maintaining low computational complexity. In summary, our contributions are as follows.
@@
    We propose an improved registration framework based on the previous work EEPNet [17] that leverages range map and reflectance map features for cross-modal registration between LiDAR point clouds and camera images, which can achieve a registration accuracy rate of over 99% in the KITTI dataset.
@@
    We introduce a patch-to-pixel feature matching framework, which not only resolves the supervision problem that arise from the limited overlap between the projection map and the camera image but also enhances registration accuracy.
II Related Work
II-A Off-line Calibration Methods
Traditional offline extrinsic calibration methods [7; 18] are typically conducted before system operation, using manually placed prominent markers to identify 2D-3D correspondences and then estimate the external parameters between sensors using the least-square method [19]. These markers, such as checkerboard, generally contain distinct features on points [20], lines [21], or planes [22]. After establishing the correspondence between 2D and 3D feature points [23; 24]. Such methods are performed before system deployment and can achieve high calibration accuracy. However, the marker placement and calibration process often require significant manpower and resources. If the sensor system’s relative poses are actively changed or vibrations or shocks during operation cause sensor position shifts, the calibration needs to be repeated, which is time-consuming and labor-intensive.
II-B Online Calibration Methods
Subsequently, many online calibration methods were proposed to optimize minor errors that occur during system operation, such as those caused by vibrations, after initial offline calibration. Existing approaches include using MLPs to estimate deviations [10; 9; 25], keypoint-based methods[26], and techniques that leverage objects [27] or segmentation [28; 29]. These real-time optimization techniques improve data alignment and enhance data fusion, leading to better performance in downstream tasks. However, these methods rely on the initial extrinsic parameters obtained from prior calibration and lose effectiveness when there are significant deviations between the actual and initial parameters. Moreover, deploying online calibration on top of complex offline calibration processes further increases costs.
II-C Cross-modal registration Methods
Cross-modal registration methods are often used in visual localization tasks [30; 31; 32], where the goal is to estimate the camera extrinsic parameters in the world coordinate system defined by large-scale point clouds. Traditionally, these methods are based on 2D visual descriptors [33; 34; 35] to achieve alignment. Later advancements have introduced cross-dimensional descriptors [11; 12; 36] designed for direct cross-modal matching. These methods generally involve extracting features from different modalities to create high-dimensional representations, using feature similarity in this space to establish correspondences. Subsequently, transformation matrices are estimated using techniques like EPnP [37].
@@ II-C Cross-modal registration Methods
Figure 3: Visualizations of the projection map, (a) Projected map by spherical projection. The red area represents the empty holes. (b) Projected map from modified projection method using LaserID. The red area represents the empty holes.
Figure 4: The detailed process of modified projection method using LaserID. The upper part of the image shows a projection map (exemplified by a reflectivity map), while the lower part displays the original point cloud being projected. Points A and B represent the 3D points captured by the uppermost and lowermost individual laser beams at the same azimuth angle, respectively. Pixels A and B correspond to the top and bottom pixels in the reflectivity map at the same horizontal coordinate.
III Proposed Method
III-A Problem Statement
@@
III-B Data Pre-processing
To get a better use of the geometric characteristics of LiDAR point cloud, and mitigate the impact of domain gap between the point cloud and the image, we project the original 3D point cloud into 2D representations by a specific projection method. Currently, there are various methods for such projection, such as spherical projection [2], cylindrical projection [40], and bird’s-eye view (BEV) projection [41]. Among them, the spherical projection can make good use of the characteristics of lateral scanning of LiDAR. However, according to the research by Wu et al. [41], direct spherical projection may result in numerous empty holes in the projected map, due to slight misalignments between the coordinate systems of individual lasers and the overall LiDAR system, as shown in the red area of Figure 3(a). These empty areas could negatively impact subsequent registration performance [42].
To address this issue, we use a modified the spherical projection by associating the vertical axis of the projection map with the LaserID of the LiDAR point cloud instead of the elevation angle, which can reduce the empty holes effectively, as illustrated in Figure 3(b). By using the radius in the spherical coordinate system as the information projected onto the map, we obtain the range map $\mathbf{I}_{Ra}\in\mathbb{N}^{H_{R}\times W_{R}}$. Similarly, projecting the reflectance intensity yields the reflectance map $\mathbf{I}_{Rf}\in\mathbb{N}^{H_{R}\times W_{R}}$. These maps, along with the camera image $\mathbf{I}_{C}\in\mathbb{N}^{H\times W\times 3}$, are fed into the subsequent network.
To explain this projection process vividly, we take the reflectance map $\mathbf{I}_{Rf}$ as an example, as shown in Figure 4. For an $N$-line LiDAR, because the horizontal axis of the projection map corresponds to the azimuth angle $\phi$ in the spherical coordinate system and the vertical axis corresponds to the LaserID, we can assume that we have two data points at the same azimuth angle $\phi$: point A in the first line (LaserID = 0) with spherical coordinates $(r_{A},\theta_{A},\phi)$ and point B in the last line (LaserID = $N-1$) with coordinates $(r_{B},\theta_{B},\phi)$, where, $r_{A}$ and $r_{B}$ are the spherical radii of points A and B, and $\theta_{A}$ and $\theta_{B}$ are their pitch angles. After this LaserID-based projection, they are mapped to the same column $u_{\phi}$, with Pixel A at $(u_{\phi},0)$ in the first row and Pixel B at $(u_{\phi},N-1)$ in the last row of the reflectance map [43].
@@ III-B Data Pre-processing
III-C Feature Extraction Network
@@
Figure 6: Detailed architecture of the patch-to-pixel matching network. The upper half depicts the patch matching network, while the lower half illustrates the pixel matching part. The gray dashed box highlights the Matching Module, which shares an identical structure in both part.
We employ convolutional neural networks (CNNs) to process the camera image, range map, and reflectance map. The camera encoder extracts features from $\mathbf{I}_{C}$, producing a high-dimensional feature at 1/32 of the original image scale. This small-scale, high-dimensional feature is then passed to the camera decoder. In the camera decoder network, the feature scale is gradually upsampled. First, after three deconvolution layers, we obtain the patch-level feature at 1/4 of the original image scale, $\mathbf{f}_{C}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times D_{\text{patch}}}$, where $D_{patch}$ is the number of patch-level feature channels. Then after two layers of deconvolution to achieve pixel-level features $\mathbf{F}_{C}\in\mathbb{R}^{H\times W\times D_{\text{pixel}}}$ at the original image scale, where $D_{pixel}$ is the number of pixel-level feature channels.
@@
III-D Patch-to-Pixel Matching Network
@@
The structure of our matching network is shown in Figure 6. Initially, we input the patch-level features $\mathbf{f}_{C}$ and $\mathbf{f}_{L}$, extracted by the feature extraction network, into a matching module to perform patch-level matching. The matching module is similar to the optimization network structure in Lightglue [44]. First, we compute the score matrix $\mathbf{S}$ between $\mathbf{f}_{C}$ and $\mathbf{f}_{L}$. Specifically, these features are flattened and then passed through a linear layer before performing matrix multiplication